Two cars leave town at the same time going in opposite directions. One of them travels 55 mph and the other 65 mph. In how many hours will they be 150 miles apart?
Note: Everytime I try to solve this problem I get 1 hour and 25 minutes, but my teacher says this is the wrong answer. It would really help me if you could show the work for this problem so I know what I'm doing wrong. Thank you!

Answer by mananth(12270)

One car speed = 55 mph
Another car speed = 65 mph
They are going away from each other.
Add up their speed
=120 mph
distance = 150 miles
Time = d/r
time = 150/120
time = 15/12
time = 5/4 hours
=1.25 hours
(0.25 hours is not 25 minutes)
(0.25 hours = 0.25 *60 = 15 minutes
1 hour 15 minutes